Tombert Decorating uses a job order cost system to collect the costs of its interior decorating business. Each client’s consultation is treated as a separate job. Overhead is applied to each job based on the number of decorator hours incurred. Listed below are data for the current year.

Estimated overhead $990,000
Actual overhead $1,073,800
Estimated decorator hours 40,000
Actual decorator hours 40,200


The company uses Operating Overhead in place of Manufacturing Overhead.

Compute the predetermined overhead rate. (Round answer to 2 decimal places, eg, 25.57%.) Predetermined overhead rate per decorator hour

Prepare the entry to apply the overhead for the year. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal place, e.g. 55.)

Account Titles and Explanation Debit Credit

Determine whether the overhead was under- or overapplied and by how much. Overhead

please include all answers and calculations. thank you
Answer #1

Predetermined overhead rate(POHR)= estimated manufacturing overhead/estimated machine hours

=990000/40000= $24.75 per machine hour

Applied overhead= POHR*actual machine hours

= 24.75*40200= $994950

Journal entry to apply overhead for the year:

Work in progress A/c Dr. 994950

To Operating Overhead A/c 994950

Note: the company uses operating overhead, hence the name of the account had been credited as "operating overhead" itself, instead of "manufacturing overhead"

Under applied by $78850(1073800-994950)